A company's overhead costs have been cut from $820,250 to $739,210. Estimate the percent of overhead that the company has cut, to the nearest percent.
A. 12%
B. 10%
C.13%
D. 15%

Answer by checkley77(12844)

($820,250-$739,210)/820,250=81.040/820,250=.098799 OR 10% ANSWER B).